bobrosssssssUwU:

Explain how to solve the equation and then check the solution. 9 = q – 4

ZombieLover:

Hi! what do you think would be the answer?

bobrosssssssUwU:

I don't know

ZombieLover:

Ok well first you Undo the subtraction of 4 by adding 4 to both sides of the equation. \[-4+4=?\]

ZombieLover:

then you do the same for \[9+4=?\]

ZombieLover:

so what are your two answers?

bobrosssssssUwU:

you Undo the subtraction of 4 by adding 4 to both sides of the equation. −4+4=?then you do the same for 9+4=? like that, I think

ZombieLover:

well yes but what are the two answers?

bobrosssssssUwU:

I think 13 and 0

ZombieLover:

Correct so your \[q=13\]

bobrosssssssUwU:

ok thank you

ZombieLover:

your welcome :) you can check by \[13-4=?\] which should be 9

bobrosssssssUwU:

ok